# Controlled Training Data Generation with Diffusion Models

### v1

@@ 3.3 Target Distribution Informed Generation
The adversarial formulation above finds prompts that reflect the failure modes of $f$. Without any information about the target distribution, improving the model on the worst-performing distributions is one of the best strategies one can do and, indeed, improves performance in some cases (see Fig. 4 and Fig. 5(a)). However, there are typically multiple failure modes of a given model and many possible distribution shifts that can occur at test-time. Adapting to all of them using only the first feedback mechanism could be inefficient when the goal is to adapt to a specific target distribution instead of improving the performance on average. Thus, we introduce the second feedback mechanism to inform the prompt optimization process of the target distribution. This only requires access to simple text descriptions (e.g., ‘fog’ to adapt to foggy images) or a small number ($\sim 100$) of unlabelled images.
@@
We implement the target-informed feedback mechanism using CLIP [57] guidance. Specifically, we assume access to either textual descriptions of the target image distribution $\{t_{j}\}$, a few unlabeled image samples $\{x_{j}\}$ or both. We then construct the corresponding text and image guidance embeddings as $e_{\mathrm{t}}=\mathrm{avg}(\{E_{\mathrm{t}}(t_{j})\})$ and $e_{\mathrm{i}}=\mathrm{avg}(\{E_{\mathrm{i}}(x_{j})\})$, where $E_{\mathrm{t}}$ and $E_{\mathrm{i}}$ denote, respectively, the CLIP text and image encoders, and $\mathrm{avg}$ stand for averaging. We then use the following guidance loss:
$\mathcal{L}_{\mathrm{CLIP}}(\tilde{x},c_{w})=\lambda_{\mathrm{t}}\mathcal{L}_{\mathrm{t}}(E_{\mathrm{t}}(c_{w}),e_{\mathrm{t}})+\lambda_{\mathrm{i}}\mathcal{L}_{\mathrm{i}}(E_{\mathrm{i}}(\tilde{x}),e_{\mathrm{i}}),$ (2)
@@
where we take $\mathcal{L}_{\mathrm{t}}$ to be $l_{2}$ norm between two embeddings and $\mathcal{L}_{\mathrm{i}}$ to be the negative cosine similarity, as we found it to perform the best. See the Appendix Sec. 7.7 for the results of this ablation. Note, that based on the available information, one can also use only one of the two guidance losses. Finally, we combine both adversarial, Eq. 1, and CLIP guidance, Eq. 2, losses to form the final objective:
$w_{\mathrm{GAP}}=\arg\min_{w}\mathbb{E}_{y}~{}\mathbb{E}_{\tilde{x}\sim g(y,c_{w})}\left[\mathcal{L}_{\mathrm{adv}}(f(\tilde{x}),y)+\mathcal{L}_{\mathrm{CLIP}}(\tilde{x},c_{w})\right]$ (3)
@@
We call the prompts that result from solving Eq. 3, Guided Adversarial Prompts
(GAP). See the Appendix Secs. 7.4.2 and 7.5.3 for further implementation details.
Figure 3: On Waterbirds, Guided Adversarial Prompts are able to generate counterfactual examples not present in the original training dataset and improves the data-efficiency over other methods. Left: We train a classification model on the original spuriously correlated dataset (see examples on the right) with varying number of extra data points generated using different types of prompts. We measure the accuracy on a balanced set where waterbirds and landbirds appear on both land and water. We run each experiment with three seeds and report the mean and standard deviation. Right: We show examples of generated images with each type of prompts on the Waterbirds. We also provide original training examples where the background is a perfectly predictive feature of the bird type. A red frame signifies that the model trained only on the original data (No Extra Data) misclassifies the image, and a green frame stands for the correct prediction. We observe that having a guidance mechanism towards the target image distribution consistently improves on top of the Agnostic Prompts baseline (“nature” prompt). Adversarial Prompts, while fooling the model, generates images that are different from the target distribution and, thus, not useful to adapt the model to it. Combining both mechanisms in Guided Adversarial Prompts leads to improved data efficiency. Unlike Guided Prompts that uses the same prompts to generate images for both classes, GAP finds prompts that generate images the model fails on, this leads to generation of waterbirds on land and landbirds on water, the combinations not present in the original training data, which are necessary data samples for the model to learn the bird predictive feature.
@@ 4 Experiments
We perform experiments in three settings: domain generalization via camera trap animal classification on the iWildCam [4] dataset, bird classification with spurious correlation with the Waterbirds [66] dataset, and depth estimation with the Taskonomy dataset [80, 79]. For depth estimation, the considered distribution shifts are Common Corruptions [29] (CC), 3D Common Corruptions [37] (3DCC) applied on the Taskonomy [80] test set and cross dataset shift from the Replica [73] dataset.
4.1 Semantic Classification
@@
Waterbirds [66] is a dataset constructed by pasting an image of either a waterbird or landbird from the CUB [75] dataset, which represents the label $y$, onto a “land” or “water” background image from the Places [84] dataset. We follow [18] and take only images of waterbirds appearing on water and landbirds on land background as $\mathcal{D}_{\mathrm{train}}$, i.e., in the training data, the background alone is predictive of the bird class. The examples from the test distribution $\mathcal{D}_{\mathrm{test}}$ contain all four combinations of bird class and background.
@@
iWildCam [3, 42] is a domain generalization dataset, made up of a large-scale collection of images captured from camera traps placed in various locations around the world. We seek to learn a model that generalizes to photos taken from new camera deployments. We follow [18] and subsample the dataset to create a 7-way classification task (background, cattle, elephant, impala, zebra, giraffe, dik-dik), with 2 test locations that are not in the training or validation set. We also fixed the number of additional generated images for finetuning to 2224 images in order to match the setting of [18].
For each dataset, we compare the following methods (we provide more details in Appendix Sec. 7):
@@
  • No Extra Data: We train a ResNet50 [26] model using the original training data $\mathcal{D}_{\mathrm{train}}$ without using any extra data.
@@
  • Augmentation baselines: We compare to two data augmentation baselines taken from recent literature: CutMix [78] and RandAugment [12].
@@
  • Agnostic Prompts: We use a prompt that is not informed of the model or the target distribution. Similar to ALIA [18], we use a prompt “nature” for Waterbirds and the prompt template “a camera trap photo of {class name}”.
@@
  • Guided Prompts (ALIA [18]): This approach uses a captioning and language model to summarize a target distribution shift into text prompts. Specifically, we use the prompts found by the ALIA method. This results in seven prompts for Waterbirds and four prompts for iWildCam. See Appendix Sec. 7.2 for more details and a discussion on the differences between ALIA and our method.
@@
  • Adversarial Prompts: We use the model trained without extra data as the target model $f$ and find adversarial prompts following Eq. 1. We find four prompts per class for Waterbirds, eight in total, and four prompts in total applied to all classes for iWildCam.
@@
  • Guided Adversarial Prompts: We use the same setting as in Adversarial Prompts and apply additional CLIP guidance to adapt to a target distribution shift, following Eq. 3. For Waterbirds, we apply text guidance using ALIA prompts as the textual description of the target distribution. For iWildCam, we use image guidance and partition the target test distribution into four groups based on two attributes that have significant impact on the visual characteristics of the data; the test location (first or second) and time of the day (day or night). We sample 64 unlabelled images randomly from each group. We optimize for one guided adversarial prompt per group.
Training details. For both datasets, we perform adversarial optimization with constant learning rate of 1e-3 using Adam [38]. We use the DDIM [72] scheduler. See Tab. 1 for a summary of the experimental parameters. The adversarial loss $\mathcal{L}_{\mathrm{adv}}$ from Eq. 1 is defined as the negative cross-entropy loss for Waterbirds. As mentioned in Sec. 3.2, we optimize the entropy loss for iWildCam. More precisely, this loss is equal to the cross entropy loss where the target label $y$ is replaced by the soft label $\tilde{y}=\{\frac{1}{|\mathcal{Y}|},\dots,\frac{1}{|\mathcal{Y}|}\}$, the uniform distribution over all classes. This loss explicitly encourages generations that either (1) do not contain new animals (2) contain new animals that are not accounted for in the label space $\mathcal{Y}$. For more details, see the Appendix.
Table 1: Adversarial optimization parameters for the classification experiments.
Waterbirds 1. 5 7.0 5 1000/1000 20/0
iWildCam 0.8 5 5.0 10 2000/10000 0/10
@@
Waterbirds results. Fig. 3-left shows the performance of each method, and Fig. 3-right demonstrates the corresponding generated examples for each method. First, we find that while Adversarial Prompts works on par with the Agnostic Prompts in a low-data regime, it performs worse with more generated data. Looking at the corresponding generations, we see that, while being adversarial, these images are different from the target distribution, which can explain the inferior performance on this particular target distribution. Second, using text prompts informed of the target leads to a consistent improvement over the Agnostic Prompts, and the corresponding images look more similar to the real images from the target distribution. Finally, Guided Adversarial Prompts combining both feedback mechanisms results in more data-efficient generations outperforming all other methods in the low-data regime. Looking at the corresponding generations, we can see that GAP tends to generate only useful combinations, i.e., waterbirds on land and landbirds on water that are missing in the training dataset. This can explain its higher sample efficiency.
Table 2: Model feedback generates data tailored to a specific model. We show test accuracy (averaged over 3 seeds) on iWildCam for different combinations of model and generated data. Each row represents the model that was used for model-based feedback to get Guided Adversarial Prompts. Each column represents the model used for finetuning on the generated data. The best dataset for each model is the one generated using that model.
Finetuned model
Data ResNet ViT
Agnostic Prompts 72.94 73.07
GAP from ResNet 83.97 72.61
GAP from ViT 83.87 77.21
iWildCam results. Fig. 4 shows that having model-informed feedback (Adversarial Prompts) helps to generate more useful data than no feedback mechanism (Agnostic Prompts) and also improves the performance of the target-only informed Guided Prompts method in the low-data regime. Guided Adversarial Prompts combines the benefits of both model- and target-informed feedback mechanisms, consistently outperforming other methods. Fig. 4 shows exemplar generations for each method. We find that while AP generates images distinct from the target distribution and images generated by target-informed methods (snow background vs. grass background), training a model using these examples in the low-data regime performs better than GP and similar to GAP.
@@
Figure 4: Left: For iWildCam, Guided Adversarial Prompts are superior in performance and data-efficiency. We train a model on the combination of the original training data and extra data generated using different types of prompts. We show the average accuracy on two iWildCam test camera trap locations. We run each experiment with three seeds and report the mean and standard deviation. (1) Guided Adversarial Prompts consistently outperforms Guided Prompts and Adversarial Prompts. For AP and GAP, there is a much lower performance drop (1-3%) when reducing the amount of generated data compared to GP (9%). This suggests that being only target-informed requires a more exhaustive exploration of the image space to find samples that are ”useful” for the pre-trained model, compared to being model-informed. Right: Qualitative results and comparison with ALIA on the iWildCam dataset. From left to right: the real training data from iWildCam; generation using the Agnostic Prompts template: “a camera trap photo of {class name}”; Guided Prompts (ALIA) generations with target-informed prompt template: “a photo of {class name} in a grassy field with trees and bushes”; Adversarial Prompts generations; and Guided Adversarial Prompts generations with the same target distribution as Guided Prompts. GAP generations, while being visually coherent with the “grassy field” location, introduce an adversarial “camouflage” effect. White arrows point to the animals for clarity. From top to bottom, they are (mis)classified as cattle, dik-dik, cattle, cattle.
Fine-tuning with Guided Adversarial Prompts from a different model. To evaluate our method’s ability to customize data to a model, $f$, and assess its resilience to changes in model architecture, we changed the model from a ResNet50 [26] to a ViT-B-16 [16] model. In Tab. 2, we investigate whether prompts optimized using one model feedback can be useful for finetuning another model, and conclude that model feedback does indeed generate data more useful for that particular model. Furthermore, the data generated from one model when used to fine-tune another model can still significantly improve performance over Agnostic Prompts. See Appendix Fig. 14 for the results of Fig. 4 with a ViT-B-16 model, similar trends hold.
@@ 4.2 Depth Estimation
For depth estimation, we consider the following pre-trained models as $f$: 1) a U-Net [63] model trained on the Taskonomy dataset [80, 79] and 2) a dense prediction transformer (DPT) [58] model trained on Omnidata [20].
We compare the following methods. They all involve fine-tuning $f$, but on different datasets. We use ControlNet v1.0 with depth conditioning for the experiments in this section. See Fig. 5 for a comparison of the generations:
@@
  • Control (No extra data): We fine-tune $f$ on the original training data. This baseline is to ensure that the difference in performance is due to the generated data, rather than e.g., longer training or optimization hyperparameters.
@@
  • Agnostic Prompts: This baseline generates data that is agnostic to the model or the target distribution. We generate images with the prompt “room” as the datasets consist of indoor images from mostly residential buildings.
  • Agnostic Prompts (Random): We generate data with “random” prompts. In our proposed method, we optimize for $n$ embedding vectors, resulting in a prompt, $c$. Thus, to match this setting, from a Gaussian distribution fitted on the embeddings from the vocabulary, we sample $n$ random embeddings to create a random prompt to be used in the data generation.
  • Adversarial Prompts: We perform the optimization as described in Eq. 1 and fine-tune on this data.
  • Guided Prompts: We perform the optimization using only the loss described in Eq. 2 and fine-tune on this data.
@@
  • Guided Adversarial Prompts: In addition to optimizing the adversarial loss, we also optimize the CLIP guidance loss as described in Eq. 3. This allows us to generate data that is also informed of a certain distribution shift.
@@
Training details. The adversarial optimization was done with AdamW [46], learning rate of $5.0\times 10^{-4}$, weight decay of $1.0\times 10^{-3}$, and batch size of 8. We set the early stopping threshold (mentioned in Sec. 3.2) to 0.08 for the UNet model and 1.0 for the DPT model. They were trained with $\ell_{1}$ and Midas loss [20] respectively. We perform a total of 30 runs to get different Adversarial Prompts. We use the DDIM [72] scheduler. During optimization, we use only 5 denoising steps, as it is more stable. For Guided Adversarial Optimization, the guidance coefficient for text and image guidance is 1 and 5 respectively. For fine-tuning, we generate images with 15 steps. For the GP runs with SDEdit, we used strength 0.6, for the GAP runs, strength 0.9. See the Appendix Sec. 8.1 for further details.
@@
Figure 5: Left: A comparison of generations with different prompts. A comparison of the generated data for the baselines and our method. Generations with Adversarial Prompts results in diverse styles that are distinct from the original training data (see fifth column). Adversarial Prompts with SDEdit results in generations shown in the fourth column. As these generations are conditioned on the original image, they look more similar to them. The last 4 columns show the generations with text guidance for target shifts fog and blur. Using Guided Prompts alone, in this case “fog” or “blur”, results in generations with a mild fog or blur. Performing Guided Adversarial Prompts results in generations with more severe fog or blur. Right: Performance of GAP with different amount of added data. The distribution shift, in this case, is defocus blur applied on the Taskonomy test set. The plot shows the $\ell_{1}$ loss (×$\times 100$) of the U-Net model versus the amount of extra data generated and used for fine-tuning. In this example, we performed image guidance with a set of 100 unlabelled corrupted RGB images. The plot shows that both GP and GAP were able to guide the optimization toward generating training data relevant to the distribution shift. Furthermore, there is a large improvement compared to using AP or the baselines with as little as 10 extra generated samples. See Appendix Fig. 18 for results on other corruptions.
(a) Quantitative results on depth estimation. $\ell_{1}$ errors on the depth prediction task for a pre-trained U-Net and DPT model. (Lower is better. U-Net losses are multiplied by 100 and DPT losses by 10, for readability). Note that these models were trained with different losses, $\ell_{1}$ for the former and Midas loss [19] for the latter, thus their performance is not comparable. We evaluate on distribution shifts from Common Corruptions (CC), 3D Common Corruptions (3DCC), and cross-datasets (CDS), Replica. The results from CC and 3DCC are averaged over all distortions and severity levels on Taskonomy. Our method is able to generate training data that can improve results over the baselines on several distribution shifts. Performing AP with SDEdit gives better results than AP under distribution shifts. Thus, conditioning on the original image seems to be helpful for these shifts. For the DPT model, the trends are similar, AP performs better than the baselines. See Appendix Tab. 6 for results on other baselines, in particular, data augmentation ones.
@@
U-Net DPT Taskonomy Replica Taskonomy Shift Clean CC 3DCC CDS CC 3DCC Control (No extra data) 2.35 4.93 4.79 5.38 3.76 3.42 Agnostic Prompts 2.47 5.03 4.17 5.30 4.06 3.58 Agnostic Prompts (Random) 2.38 4.96 4.11 5.14 3.88 3.51 Adversarial Prompts 2.49 4.36 4.02 5.12 3.40 3.28 Adversarial Prompts (SDEdit) 2.59 4.20 3.88 4.96 3.35 3.25
(b) Comparing the performance from running multiple iterations versus a single iteration of adversarial optimization, generation, and fine-tuning. The plot shows the $\ell_{1}$ loss (×$\times 100$) of the U-Net model against the number of iterations on the depth prediction task. The $\ell_{1}$ loss is computed on the Taskonomy dataset under common corruptions, averaged over all corruptions and severity levels. The single- and multi-iteration runs have similar settings in terms of the total number of AP used, the same number of tokens per prompt, etc. See Appendix Sec. 8.4.1 for further details. The first iteration of the multi-iteration run resulted in the largest improvement in performance. However, it converges to a similar performance as the single iteration run. Thus, we chose to perform only a single iteration in Fig. 5(a).
Figure 6: Left: Results for Adversarial Prompts against the baselines. Right: Results from performing multiple iterations of adversarial optimization and finetuning.
@@
Comparing the generated images with different prompts. Fig. 5-left shows the results of the generations for the baselines and our method, optimized on the Taskonomy dataset. The generation with Agnostic Prompts are visually different from that of the original image, however, they tend to have similar styles. In contrast, the generations with Adversarial Prompts have more complex styles and are more diverse. Using SDEdit during the optimization and generation results in generations that are closer to the original image, as it was also used as conditioning. The last four columns show the results of using CLIP text guidance for the target distribution shift fog and blur, as described in Sec. 3.3, with and without adversarial optimization. The generations with Guided Prompts involve passing the depth conditioning and prompt “fog” or “blur” to the diffusion model. In both cases, the generations result in a mild level of fog or blur. In contrast, Guided Adversarial Prompts results in more severe fog and blur corruptions. See Appendix Sec. 8.4.2 for generations using image guidance.
Note that all the generations follow the conditioning, i.e., depth labels (see first column). See Sec. 3.2 for the discussion on how we prevent the generations from collapsing. Thus, this gives us aligned training data, i.e., RGB images and depth labels that we can use for fine-tuning.
Performance on OOD data. We evaluate our method and the baselines after fine-tuning on their respective generated datasets. Tab. 5(a) demonstrated that data generated with Adversarial Prompts improve the performance of the model under different distribution shifts. Furthermore, we show that the trend holds with the DPT model. Thus, our method is able to successfully find useful Adversarial Prompts for different architectures.

### v2

@@ 3.3 Target Distribution Informed Generation
The adversarial formulation above finds prompts that reflect the failure modes of $f$. Without any information about the target distribution, improving the model on the worst-performing distributions is one of the best strategies, which, indeed, improves performance in multiple cases (see Fig. 4 and Tab. 1). However, a given model typically has many failure modes and many possible distribution shifts that can occur at test time. Adapting to all of them using only the adversarial feedback mechanism might be inefficient if the goal is to adapt to a specific target distribution instead of improving average performance. Thus, we introduce the second feedback mechanism to inform the prompt optimization process of the target image distribution. It only requires access to simple text descriptions (e.g., ‘fog’ to adapt to foggy images) or a small number ($\sim 100$) of unlabelled images.
@@
We implement the target-informed feedback mechanism using CLIP (Radford et al., 2021) guidance. Specifically, we assume access to either textual descriptions of the target image distribution $\{t_{j}\}$, a few unlabeled image samples $\{x_{j}\}$, or both. We construct the corresponding text and image guidance embeddings as $e_{\mathrm{t}}=\mathrm{avg}(\{E_{\mathrm{t}}(t_{j})\})$ and $e_{\mathrm{i}}=\mathrm{avg}(\{E_{\mathrm{i}}(x_{j})\})$, where $E_{\mathrm{t}}$ and $E_{\mathrm{i}}$ denote, respectively, the CLIP text and image encoders, and $\mathrm{avg}$ stand for averaging. We then use the following guidance loss:
$\mathcal{L}_{\mathrm{G}}(\tilde{x},c_{w})=\lambda_{\mathrm{t}}\mathcal{L}_{\mathrm{t}}(E_{\mathrm{t}}(c_{w}),e_{\mathrm{t}})+\lambda_{\mathrm{i}}\mathcal{L}_{\mathrm{i}}(E_{\mathrm{i}}(\tilde{x}),e_{\mathrm{i}}),$ (2)
@@
where we take $\mathcal{L}_{\mathrm{t}}$ to be $l_{2}$ norm between two embeddings and $\mathcal{L}_{\mathrm{i}}$ to be the negative cosine similarity, as we found it to perform the best. See the Sec. B.7 for the results of this ablation. Note that our formulation allows the use of only text or only images for guidance (setting corresponding $\lambda$s to 0) based on the available information. In Sec. B.6, we also show that one can construct a similarly effective guidance mechanism without relying on the CLIP model.
Finally, we combine both adversarial, Eq. 1, and guidance, Eq. 2, losses to form the final objective:
$w_{\mathrm{GAP}}=\arg\min_{w}\mathbb{E}_{y}~{}\mathbb{E}_{\tilde{x}\sim g(y,c_{w})}\left[\mathcal{L}_{\mathrm{adv}}(f(\tilde{x}),y)+\mathcal{L}_{\mathrm{G}}(\tilde{x},c_{w})\right].$ (3)
@@
We call the prompts that result from solving Eq. 3, Guided Adversarial Prompts (GAP). See the Secs. B.4.2 and B.5.3 for further implementation details.
@@ 4 Experiments
We perform experiments in three settings: domain generalization via camera trap animal classification on the iWildCam (Beery et al., 2021) dataset, bird classification with spurious correlation with the Waterbirds (Sagawa et al., 2019) dataset, and depth estimation with the Taskonomy dataset (Zamir et al., 2018b; 2020). For depth estimation, we evaluate on distribution shifts from Common Corruptions (Hendrycks & Dietterich, 2019) (CC), 3D Common Corruptions (Kar et al., 2022) (3DCC) applied on the Taskonomy (Zamir et al., 2018b) test set and cross dataset shift from the Replica (Straub et al., 2019) dataset.
4.1 Semantic Image Classification
@@
Waterbirds (Sagawa et al., 2019) is a dataset constructed by pasting an image of a waterbird or landbird from the CUB (Wah et al., 2011) dataset, which represents the label $y$, onto a background image of “land” or “water” from the Places (Zhou et al., 2014) dataset. We follow Dunlap et al. (2023) and use only images of waterbirds on water and landbirds on land as $\mathcal{D}_{\mathrm{train}}$, i.e., the background is completely correlated with the label in training data (see Fig. 3-right). The test set $\mathcal{D}_{\mathrm{test}}$ contains all four combinations of bird types and backgrounds (see Fig. 7).
@@
iWildCam (Beery et al., 2021; Koh et al., 2021) is a domain generalization dataset made up of images captured from camera traps placed in various locations around the world. The goal is to generalize to photos taken from new camera deployments. We follow Dunlap et al. (2023) and create a 7-way classification task (background, cattle, elephant, impala, zebra, giraffe, dik-dik), use two locations that are not present in the training or validation as $\mathcal{D}_{\mathrm{test}}$, and fix the number of additional generated images for finetuning to 2224.
For both datasets, we use the ResNet50 (He et al., 2016a) model. We compare the following methods for generating additional synthetic data $\mathcal{D}_{\mathrm{syn}}$ (we provide more details in Appendix B):
@@
  • No Extra Data: Only $\mathcal{D}_{\mathrm{train}}$ is used, without generating extra data.
@@
  • Augmentation baselines: We use two data augmentation methods commonly used in recent literature: CutMix (Yun et al., 2019) and RandAugment (Cubuk et al., 2020). These baselines do not use generative models.
@@
  • Agnostic Prompts: We use a prompt that is uninformed of both the model and the specific target distribution. Similar to ALIA (Dunlap et al., 2023), we use a prompt “nature” for Waterbirds and the prompt template “a camera trap photo of {class name}” for iWildCam.
@@
  • Guided Prompts: We compare to ALIA (Dunlap et al., 2023), which uses captioning and language models to summarize a target distribution shift into text prompts. This results in seven prompts for Waterbirds and four prompts for iWildCam. See Sec. B.2 for more details, specific prompts, and a discussion on the differences between ALIA and our method.
@@
  • Adversarial Prompts: We use the model previously trained on $\mathcal{D}_{\mathrm{train}}$ (i.e., no extra data) as the target model $f$ and find adversarial prompts following Eq. 1. We find four prompts per class for Waterbirds, eight in total, and four prompts in total applied to all classes for iWildCam.
@@
  • Guided Adversarial Prompts: We use the same setting as in Adversarial Prompts and apply additional CLIP guidance to adapt to a target distribution shift, following Eq. 3. For Waterbirds, we apply text guidance using ALIA prompts as the textual description of the target distribution. For iWildCam, we use image guidance and partition the target test distribution into four groups based on two attributes that have a significant impact on the visual characteristics of the data: the test location (first or second) and time of the day (day or night). We sample 64 unlabelled images randomly from each group. We optimize for one guided adversarial prompt per group, resulting in four prompts in total.
Training details. For supervised training, we follow Dunlap et al. (2023) and use the ResNet50 model for $f$. In Sec. 4.3, we also explore how adversarial prompts change with the model by using ViT (Dosovitskiy et al., 2020). For prompt optimization, we use Adam (Kingma & Ba, 2017) with 1e-3 learning rate. For efficiency, we use five denoising steps during optimization and 15 for the final generations, both using the DDIM (Song et al., 2020) scheduler. As $\mathcal{L}_{\mathrm{adv}}$, we use prediction’s entropy for iWildCam to avoid alignment collapse between $\tilde{x}$ and $y$ as discussed in Sec. 3.2, and negative cross-entropy for Waterbirds where we did not observe collapse. Please see Appendix B for more details on the experimental setup.
Figure 3: Guided Adversarial Prompts generate counterfactual examples not present in the original training dataset and achieve better data efficiency than other prompts. Left: Test performance on a balanced test set (waterbirds and landbirds appear on both land and water) of a model trained on the original spuriously correlated dataset, while varying the number of extra data points generated by different types of prompts (see Sec. C.2.3 for results on 2X data). We run each experiment with three seeds and report the mean and standard deviation. Right: Exemplary generations for each prompt type and original training examples (where background is correlated with the bird type.) Frames show whether the model trained without extra data made a correct prediction (green) or not (red) on the corresponding image. As expected, the adversarial mechanism produces examples that fool the model, and the guidance mechanism steers the generation toward the target distribution. Combining both mechanisms in GAP leads to the generation of counterfactual examples (waterbirds on land and landbirds on water), missing in the training dataset, and improved data efficiency compared to only target-informed GP.
@@
Waterbirds results. Fig. 3 (Left) shows the performance of each method, and Fig. 3 (Right) demonstrates the corresponding generated examples for each method. First, we find that while the performance of Adversarial Prompts is on par with Agnostic Prompts in a low-data regime, it performs worse with more generated data. Looking at the corresponding generated images, we find that while being adversarial, they appear different from images from the target distribution, which can explain the inferior performance in this particular case. Second, using Guided Prompts informed of the target distribution leads to a consistent improvement over Agnostic Prompts, and the corresponding images look more similar to the real images from the target distribution. Finally, combining both feedback mechanisms in Guided Adversarial Prompts results in more data-efficient generations, outperforming all other methods in the low-data regime. Looking at the corresponding generations, we find that GAP tends to generate only combinations of the bird type and background that are missing in the training dataset, i.e., waterbirds on land and landbirds on water, which can explain its higher sample efficiency.
iWildCam results. Fig. 4 shows that similar to Waterbirds, GAP, combining both feedbacks, achieves the best performance. Compared to GP with the target-only feedback, GAP generates images with a “camouflage” effect that fools the classification model. Compared to AP with the model-only feedback that generates images with snow background, GAP’s images are more similar to the target domain that is summarized as “grassy field” (by GPT-4 in (Dunlap et al., 2023)). We also find that in the case of iWildCam, AP significantly outperforms GP in the low-data regime, suggesting that model-informed feedback generates more useful training examples. We also find that AP and GAP outperform standard (non-generative) augmentation baselines and domain generalization baselines (see Sec. C.2.2 for results).
@@
Figure 4: Guided Adversarial Prompts generate hard training examples leading to the best performance. Left: Performance on the images from unseen camera locations (out-of-distribution) with different numbers of generated synthetic images for iWildCam. We repeat each experiment three times and report the mean and standard deviation. GAP consistently outperforms other types of prompts. We also find that AP significantly outperforms GP in the low-data regime, suggesting that being only target-informed requires a more exhaustive sampling to find useful training examples. Right: Original training examples from iWildCam and generated examples for each type of prompt. For GAP, arrows point to the animals for clarity; from top to bottom, they are misclassified as cattle, dik-dik, cattle. GAP generates images that are both visually coherent with the target location (described as “grassy field”) and introduce an adversarial “camouflage” effect.
@@ 4.2 Depth Estimation
We use the Taskonomy (Zamir et al., 2018a) and Omnidata Eftekhar et al. (2021a) datasets for training. The former consists of 4 million images of indoor scenes, while the latter is a mixture of datasets and contains both indoor and outdoor scenes. We evaluate our method on a range of domain shifts, from Common Corruptions (Hendrycks & Dietterich, 2019), 3D Common corruptions (Kar et al., 2022) to other datasets like Replica. We compare the following methods. They all involve fine-tuning the originally trained model $f$ on different synthetically generated datasets. See Fig. 5 for a comparison of the generations:
@@
  • Control (No extra data): We fine-tune $f$ on the original training data. This baseline is to ensure that the difference in performance is due to the generated data rather than, e.g., longer training or other hyperparameters used during fine-tuning.
@@
  • Agnostic Prompts: This baseline generates data that is agnostic to the model and the target distribution. We generate images with the prompt “room” as the datasets consist of indoor images from mostly residential buildings.
  • Agnostic Prompts (Random): As we use 30 tokens for our methods, this baseline controls for the number of tokens used. Here, we sample the same number of tokens as AP randomly from the vocabulary to construct a prompt. See Sec. C.1 for details.
  • Adversarial Prompts: We perform the optimization as described in Eq. 1, to get 30 adversarial prompts (tokens).
  • Guided Prompts: We generate data using only the CLIP guidance loss described in Eq. 2. We consider both image and text guidnace. For the former we use about 100 unlabelled images from the target distribution, for the latter we use the name of the shift e.g., “fog”.
@@
  • Guided Adversarial Prompts: We combine adversarial and guidance losses. We use the same settings as Adversarial Prompts and apply CLIP guidance loss used for Guided Prompts. This allows us to generate data that is both useful for the model and target distribution.
@@
Training details. We consider the following pre-trained models for $f$: 1) a U-Net (Ronneberger et al., 2015) model trained on the Taskonomy dataset (Zamir et al., 2018b; 2020) and 2) a dense prediction transformer (DPT) (Ranftl et al., 2021) model trained on Omnidata (Eftekhar et al., 2021b). The adversarial optimization was done with AdamW (Loshchilov & Hutter, 2019), learning rate of $5.0\times 10^{-4}$, weight decay of $1.0\times 10^{-3}$. We set the early stopping threshold (mentioned in Sec. 3.2) to 0.08 for the UNet model and 1.0 for the DPT model. They were trained with $\ell_{1}$ and Midas loss (Eftekhar et al., 2021b) respectively. We perform a total of 30 runs to get 30 different Adversarial Prompts. Similar to classification, for efficiency, we use 5 denoising steps during optimization and 15 for the final generations, with the DDIM (Song et al., 2020) scheduler, which we found to be a good trade-off between quality and efficiency. See the Sec. C.1 for more details.
@@
Figure 5: Left: A comparison of generations with different prompts. Generations with AP results in diverse styles that are distinct from the original training data and with agnostic prompts. AP (SDEdit) generations have slight modifications to the original image, as it was conditioned on them. The last 4 columns show the generations with text guidance for target shifts fog and blur. Compared to GP, GAP results in generations with more severe fog or blur. Right: Performance of GAP with different amount of added data. The target distribution here is defocus blur applied to Taskonomy test images. The plot shows the $\ell_{1}$ loss of the UNet model versus the amount of extra data generated and used for fine-tuning. The plot shows that both GP and GAP were able to guide the optimization toward generating training data relevant to the distribution shift. See Fig. 18 for results on other corruptions and Sec. C.2.3 for results on $10^{5}$ data.
Comparing the generated images with different prompts.
Table 1: Quantitative results for depth estimation. $\ell_{1}$ errors for the depth prediction task for U-Net and DPT models on Common Corruptions (CC), 3D Common Corruptions (3DCC), and cross-datasets (CDS) like Replica. Results for CC and 3DCC are averaged over all distortions and severity levels. AP outperform baselines for all models and distribution shifts. As GAP requires optimizing for each shift individually, and the benchmarks contain $\sim 30$ shifts in total, we only perform GAP for a few shifts (see Fig. 5 and Fig. 18 for results). See Tab. 7 for results on other baselines.
@@
UNet (×100)($\times$ 100$\downarrow$)( × 100 ↓ ) DPT (×10)($\times$ 10$\downarrow$)( × 10 ↓ ) Taskonomy Replica Taskonomy Shift Clean CC 3DCC CDS CC 3DCC Control (No extra data) 2.35 4.93 4.79 5.38 3.76 3.42 Agnostic Prompts 2.47 5.03 4.17 5.30 4.06 3.58 Agnostic Prompts (Random) 2.38 4.96 4.11 5.14 3.88 3.51 Adversarial Prompts 2.49 4.36 4.02 5.12 3.40 3.28 Adversarial Prompts (SDEdit) 2.59 4.20 3.88 4.96 3.35 3.25
@@
Fig. 5-left shows the generations from the baselines and our methods. The generations with AP are diverse and have complex styles. Incorporating SDEdit results in slight modifications to the original image, as it was used for conditioning. The last four columns show the results of using text guidance for the target distribution shift fog and blur (see Sec. C.4.2 for a comparison between generations from image and text guidance). The generations for GP results only in a mild level of fog or blur. In contrast, GAP results in more severe fog and blur corruptions. The generations for GP, using only the guidance feedback, result only in a mild level of fog or blur. Finally, incorporating both feedback mechanisms in GAP results in generating images with corruptions that are both relevant to the corresponding target distribution and more severe, making them more useful for improving the model.